3 June 2014: Tuesday

LDAP:

  • No issues in last night’s update.
  • Dealing with FP ISSUE=2698 PROJ=128 – a student dropped out of the Banner feed to LDAP and doesn’t think they should have.  Something happened to the Banner feed that they dropped out.  The person who needs to examine it will be back next week.  As an interim fix, we restored the student’s data in LDAP from April 29.

BACKUPS:

  • Those two new clients didn’t work the way they should have.  One of them appears to have a closed firewall (or is not running) the other is not backing up the filesystems (linux ZFS is not a supported system anyway).
    • Firewall issue on the one that completely failed – fixed.
    • Yeah, linux ZFS is going to present a fun issue.
  • Tape clones finished last night, Disk and CoM/IS clones continue to run.
  • 43 tapes should have recycled themselves already, they have not… manually seeing why and forcing the ones that should go.  All forty-three (43) were found wanting and relabeled.

Tags:

Leave a Reply

You must be logged in to post a comment.